### v3

# A tweezer array with 6100 highly coherent atomic qubits

Abstract
Optical tweezer arrays browaeys_many-body_2020, kaufman_quantum_2021 have had a transformative impact on atomic and molecular physics over the past years, and they now form the backbone for a wide range of leading experiments in quantum computing saffman_quantum_2016, henriet_quantum_2020, bluvstein_quantum_2022, graham_multi-qubit_2022, ma_universal_2022, finkelstein_universal_2024, simulation bernien_probing_2017, browaeys_many-body_2020, ebadi_quantum_2021, scholl_quantum_2021, shaw_benchmarking_2024, and metrology norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020. Underlying this development is the simplicity of single-particle control and detection inherent to the technique barredo_atom-by-atom_2016, endres_atom-by-atom_2016, kim_situ_2016, liu_building_2018, anderegg_optical_2019. Typical experiments trap tens to hundreds of atomic qubits young_half-minute-scale_2020, scholl_quantum_2021, huft_simple_2022, singh_mid-circuit_2023, tao_high-fidelity_2024, bluvstein_logical_2024, shaw_benchmarking_2024, and very recently systems with around one thousand atoms were realized without defining qubits or demonstrating coherent control pause_supercharged_2024, norcia_iterative_2024_mod, gyger_continuous_2024. However, scaling to thousands of atomic qubits with long coherence times and low-loss, high-fidelity imaging is an outstanding challenge and critical for progress in quantum computing, simulation, and metrology, in particular, towards applications with quantum error correction preskill_quantum_2018, xu_constant-overhead_2024. Here, we experimentally realize an array of optical tweezers trapping over 6,100 neutral atoms in around 12,000 sites while simultaneously surpassing state-of-the-art performance for several critical metrics that underpin the success of the platform. Specifically, while scaling to such a large number of atoms, we also demonstrate a coherence time of 12.6(1) seconds, a record for hyperfine qubits in an optical tweezer array bluvstein_quantum_2022, graham_multi-qubit_2022. Further, we show trapping lifetimes close to 23 minutes in a room-temperature apparatus, enabling record-high imaging survival covey_2000-times_2019, blodgett_imaging_parsed_2023 of in combination with an imaging fidelity of over . We lay out a detailed, near-term plan to enable zone-based quantum computing bluvstein_quantum_2022, bluvstein_logical_2024 with atoms, and demonstrate a crucial ingredient thereof: coherent moves of up to m with a fidelity of , as characterized through interleaved randomized benchmarking. Our results, together with other recent developments evered_high-fidelity_2023, ma_high-fidelity_2023, zhang_scaled_2024, finkelstein_universal_2024, indicate that universal quantum computing bluvstein_quantum_2022, graham_multi-qubit_2022, ma_universal_2022, finkelstein_universal_2024 with ten thousand atomic qubits could be a near-term prospect. Furthermore, our work could pave the way for quantum simulation orourke_entanglement_2023, julia-farre_amorphous_2024 and metrology ludlow_optical_2015, rosenband_exponential_2013, haferkamp_closing_2020 experiments with inherent single particle readout and positioning capabilities at a similar scale.
Optical tweezer arrays browaeys_many-body_2020, kaufman_quantum_2021 have transformed atomic and molecular physics experiments by simplifying detection and enabling control at the individual-particle level barredo_atom-by-atom_2016, endres_atom-by-atom_2016, kim_situ_2016, liu_building_2018, anderegg_optical_2019, resulting in rapid, recent progress in quantum computing saffman_quantum_2016, henriet_quantum_2020, bluvstein_quantum_2022, graham_multi-qubit_2022, ma_universal_2022, finkelstein_universal_2024, quantum simulation bernien_probing_2017, browaeys_many-body_2020, ebadi_quantum_2021, scholl_quantum_2021, shaw_benchmarking_2024, and optical frequency metrology norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020. They are part of broader efforts towards single atom control in optical traps, including lattices miroshnychenko_atom-sorting_2006, weitenberg_single-spin_2011 and lattice-tweezer hybrid systems young_half-minute-scale_2020, tao_high-fidelity_2024, gyger_continuous_2024. For quantum computers, simulators of spin systems and tweezer clocks, each atom typically encodes a single qubit that is controlled with electromagnetic fields, and ideally features long coherence times to enable these applications with high fidelity. Such optically trapped atomic qubits coexist with other platforms that have single-qubit control and readout, including ion traps bruzewicz_trapped-ion_2019 and superconducting qubits kjaergaard_superconducting_2020.
There are important incentives to scale up such fully programmable qubit platforms. For example, optical clocks gain in stability with increasing atom number ludlow_optical_2015, rosenband_exponential_2013, and proposals have been made for quantum simulation experiments that benefit from several thousand qubits to explore emergent collective behavior orourke_entanglement_2023, julia-farre_amorphous_2024 or to demonstrate verifiable quantum advantage haferkamp_closing_2020, ringbauer_verifiable_2023.
In particular, quantum error correction (QEC) requires large system sizes, while maintaining high fidelities and long coherence times, in order to scale up the number of logical qubits preskill_quantum_2018, google_quantum_ai_suppressing_2023. Even the most resource-efficient QEC protocols require at least several thousand physical qubits with an error-per-gate level of to encode 100 error-corrected logical qubits bravyi_high-threshold_2024, xu_constant-overhead_2024. However, published results from universal quantum computing architectures based on ion traps and superconducting qubits currently leverage only tens to hundreds of qubits moses_race_2023, kim_evidence_2023, and with current technologies for fabrication and control, generally suffer from deleterious effects as they are scaled kjaergaard_superconducting_2020, bruzewicz_trapped-ion_2019.
Neutral atoms in optical tweezer arrays are a promising platform towards rapid scalability in the near term thanks to a programmable architecture that is readily adaptable to larger system sizes. In particular, universal quantum computing has recently been realized in such systems bluvstein_quantum_2022, graham_multi-qubit_2022, ma_universal_2022, finkelstein_universal_2024, based on demonstrations of individual qubit addressing huie_repetitive_2023, lis_midcircuit_2023, norcia_midcircuit_2023, shaw_multi-ensemble_2024, high-fidelity entangling gates evered_high-fidelity_2023, finkelstein_universal_2024, and coherence-preserving dynamical reconfigurability beugnon_two-dimensional_2007, bluvstein_quantum_2022, finkelstein_universal_2024, alongside ancilla-based mid-circuit measurement singh_mid-circuit_2023, bluvstein_logical_2024, finkelstein_universal_2024 and erasure error detection scholl_erasure_2023, ma_high-fidelity_2023. Leveraging the versatility of the optical tweezer architecture, several applications have been shown, including executing quantum phase estimation graham_multi-qubit_2022, realizing the toric code structure bluvstein_quantum_2022, building an optical clock with ancilla-based readout finkelstein_universal_2024, and implementing logical qubit operations bluvstein_logical_2024.
In terms of current system sizes, tens to hundreds of qubits are controlled by typical tweezer array experiments young_half-minute-scale_2020, scholl_quantum_2021, huft_simple_2022, singh_mid-circuit_2023, tao_high-fidelity_2024, bluvstein_logical_2024, shaw_benchmarking_2024. Very recently, tweezer systems with about a thousand atoms have been realized in a discontiguous array based on interleaved microlens elements pause_supercharged_2024, and via repeated reloading from a reservoir norcia_iterative_2024_mod, gyger_continuous_2024, following earlier iterative reloading schemes singh_dual-element_2022, shaw_dark-state_2023; none of these experiments, however, report control of qubits or measurement of coherence times.
Here, we demonstrate a tweezer array with 11,998 sites, that traps over 6,100 atomic qubits, simultaneously surpassing state-of-the-art values for metrics underpinning the usefulness of the platform, including hyperfine qubit coherence time, trapping lifetime in a room-temperature apparatus, coherent transport distance and fidelity, as well as combined imaging fidelity and survival (Fig. 1). Our results have implications for aforementioned applications in quantum science, in particular, concerning large-scale quantum computing and error correction, as discussed in more detail below.
Summary of approach and results
Our approach leverages high-power trapping of single atoms at far-off-resonant wavelengths in a specially designed, room-temperature vacuum chamber (Methods, Ext. Data Fig. 1a), enabling low-loss, high-fidelity imaging in combination with long hyperfine coherence times at the scale of 6,100 qubits (Fig. 1e). Specifically, we demonstrate imaging of single cesium-133 atoms with a survival probability of , in combination with an imaging fidelity of , surpassing the state-of-the-art achieved in much smaller arrayscovey_2000-times_2019. This, alongside a 22.9(1) minute vacuum-limited lifetime in our room-temperature apparatus schymik_single_2021 — significantly longer than typical state-of-the-art vacuum lifetimes for tweezer arrays in room-temperature apparatuses— provides realistic timescales for array operations in large scale arrays with minimal loss, e.g., for atomic rearrangement barredo_atom-by-atom_2016, endres_atom-by-atom_2016, kim_situ_2016.
Importantly, we further demonstrate a coherence time of 12.6(1) s, a record for a hyperfine qubit tweezer array, surpassing previous values by almost an order of magnitude bluvstein_quantum_2022, graham_multi-qubit_2022 and approaching results for a single hyperfine qubit in a customized blue-detuned trap tian_extending_2024, alkali atoms in an optical lattice wu_sterngerlach_2019, and nuclear qubits in a tweezer array barnes_assembly_2022. We also show a single-qubit gate fidelity of measured with global randomized benchmarking, limited by technical noise. We observe high uniformity across the sites in the array for loading probability (Ext. Data Fig. 2c), trap depth (Ext. Data Fig. 2d), imaging survival (Ext. Data Fig. 5a), scattering rate (Ext. Data Fig. 4d), and qubit frequency (Ext. Data Fig. 8a).
Finally, we lay out a realistic plan to implement zone-based universal quantum computation bluvstein_quantum_2022, bluvstein_logical_2024 with 6,000 atoms in the near term. We summarize in this work the key elements of our proposal, elucidated further in the Supplementary Information (SI). We notably demonstrate coherent atomic transport on a distance of 610 m with a fidelity of 99.95%, a crucial ingredient for scaling this architecture. Our results indicate that quantum computing with 6,000 atomic qubits is a near-term prospect, providing a path towards QEC with hundreds of logical qubits xu_constant-overhead_2024.
Large-scale optical tweezer generation
We scale the optical tweezer array platform to 11,998 tweezers, while prioritizing achieving long qubit coherence times and high fidelities (Fig. 1e). As such, we trap using near-infrared wavelengths, far-detuned from dominant electric-dipole transitions to minimize hyperfine qubit decoherence from photon scattering and dephasing processes ozeri_hyperfine_2005, ozeri_errors_2007. Cesium atoms possess the highest polarizability among the stable alkali-metal atoms at near-infrared wavelengths where commercial fiber amplifiers routinely provide continuous-wave laser powers that exceed 100 W. Thus, a large number of traps can be generated with sufficient depth. A representative single shot image of the array is shown in Fig. 1a, and an averaged image is shown in Fig. 1b.
The atoms are spaced by and held in traps at 1055 nm and 1061 nm, generated using spatial light modulators (SLMs), whose hologram phases are optimized with a weighted Gerchberg-Saxton algorithm kim_gerchberg-saxton_2019, kim_large-scale_2019 to uniformize the tweezer trap depth across the array (Methods). The tweezer light is combined with polarization and focused through a high numerical aperture objective with an unusually large field of view of 1.5 mm diameter that provides a large area for atom trapping and manipulation in ultra-high vacuum (Fig. 1c). We measure an average trap depth of mK, with a standard deviation of across all sites (Ext. Data Fig. 2d), enabling consistent loading probability per site.
Using Raman sideband spectroscopy with a 9.2 GHz amplitude-modulation technique levine_dispersive_2022, we measure radial and axial trapping frequencies of 29.3 kHz and 5.6 kHz, respectively (for the traps generated with the 1061 nm laser only; see Ext. Data Fig. 9). The setup used to address the atomic motion with a Raman configuration is detailed in the Methods. From this measurement we infer a tweezer waist .
Loading and imaging single atoms
We demonstrate uniform loading and high imaging fidelity across the sites in the array. To load single atoms in the tweezers, we cool and then parity-project schlosser_sub-poissonian_2001 from a mm diameter magneto-optical trap. Before imaging the atoms, we use a multi-pronged approach to filter out atoms in spurious off-plane traps, residual from the SLM tweezer creation (Methods).
We then zero the magnetic field and apply two-dimensional polarization gradient cooling (2D PGC) beams that are approximately parallel to the atom array plane (- plane in Fig. 1c) for fluorescence imaging of single atoms, which also simultaneously cool the atoms. The photons scattered from each atom are collected on a quantitative CMOS (qCMOS) camera for 80 ms, using the same objective with which the traps are created. We find that each site is loaded with a single atom on average 51.2% of the time with a relative standard deviation of 3.4% across the sites, demonstrating uniform filling of single atoms in a large tweezer array (Ext. Data Fig. 2c). This allows us to load over 6,100 sites on average in each iteration (Fig. 1d).
We then detect and distinguish atomic presence in the array with high fidelity. Each image undergoes a binarization procedure (detailed in Methods) whereby each site is attributed a value of 0 (no atom detected) or 1 (one atom detected). Note that we detect no more than one atom schlosser_sub-poissonian_2001 in each tweezer. We weight the collected photons in a -pixel box centered around each site cooper_alkaline-earth_2018, so as to add more weight to pixels close to the center of each site’s point-spread function (Ext. Data Fig. 4a). The resulting signal is compared with a threshold to determine if an atom is present or not (Fig. 2).
We characterize the imaging fidelity, defined as the probability of correctly labeling atomic presence in a site, with a model-free approach, where no assumption is made about the photon distribution from Fig. 2. To this end, we identify anomalous series of binary outputs norcia_microscopic_2018 in three consecutive images. For instance, would point to a false negative event in the first image, while could be due to atom loss during the second image or a false negative event in the third one. This approach allows us to precisely decouple inherent atom loss from false negatives or positives. From this we find an imaging fidelity of 99.99374(8)% (note that we excise the first image, which we find has slightly lower fidelity and survival probability; Methods). Crucial to this result are the homogeneous photon scattering rate across the array (Ext. Data Fig. 4d) and the consistency of the point-spread function even for sites at the edge of the array (waist radius of 1.7 pixels with a standard deviation of 0.2 pixels). Furthermore, we find that treating each site separately with an individual threshold only marginally improves the imaging fidelity to 99.9939(1)%, indicating that the imaging parameters are sufficiently consistent across the atom array. Finally, we estimate that the imaging fidelity in the absence of atomic loss would be closer to 99.999% (Methods).
Imaging survival and vacuum-limited lifetime
The probability of not losing a single atom in a tweezer array during imaging and due to finite vacuum lifetime both decrease exponentially in the number of atoms in an array, making these crucial metrics to optimize for large-scale array operation. The vacuum-limited lifetime, in particular, sets the ultimate upper bound on the duration during which operations can be executed without loss of an atom in a given experimental setup. This can, for example, be applied to give an upper limit on the fidelity with which one can achieve a defect-free array via atom rearrangement barredo_atom-by-atom_2016, endres_atom-by-atom_2016.
We probe the vacuum-limited lifetime by using an empirically optimized cooling sequence consisting of a 10 ms 2D PGC cooling block every 2 s. By fitting the exponential decay of the atom survival, we find a lifetime of 22.9(1) min (Fig. 3a). This is a significantly long timescale compared with state-of-the-art room-temperature atomic experiments, and only within a factor of five of the longest reported lifetime in a cryogenic apparatus schymik_single_2021. The result indicates that the probability of losing a single atom across the entire array remains under 50% during 100 ms, a relevant timescale for dynamical array reconfiguration and quantum processor operation.
Moreover, we accurately characterize the imaging survival probability, without assuming any parameters, by performing 80 ms repeated imaging up to 1,000 times, after which % of initially loaded atoms still survive (Fig. 3b). This corresponds to a steady-state imaging survival probability of , mostly limited by vacuum lifetime. This, to the best of our knowledge, surpasses prior studies reporting record steady-state imaging survival using single alkaline-earth metal covey_2000-times_2019 and alkali-metal blodgett_imaging_parsed_2023 atoms in optical tweezers. These results, and the uniformity of imaging survival across the array (Ext. Data Fig. 5a), enable low-loss, high-fidelity detection of single atoms in large-scale arrays, crucial components for the practical use of the system.
Qubit coherence
Key to recent progress in quantum computing and metrology with neutral atoms is the ability to encode a qubit in long-lived states of an atom, such as hyperfine states bluvstein_quantum_2022, graham_multi-qubit_2022, nuclear spin states barnes_assembly_2022, huie_repetitive_2023, lis_midcircuit_2023, norcia_midcircuit_2023, or optical clock states norcia_seconds-scale_2019, madjarov_atomic-array_2019. In cesium atoms, the hyperfine ground states ( and ) provide such a subspace for storing quantum information (see Methods for state preparation and readout procedures). Furthermore, entanglement via Rydberg interactions can be readily transferred to this qubit to realize high-fidelity two-qubit gates evered_high-fidelity_2023, finkelstein_universal_2024. We now demonstrate the storage and manipulation of quantum information in a large-scale atom array by measuring the coherence time and global single-qubit gate fidelity. To this end we use a microwave horn to drive the hyperfine transition (Fig. 4a). For microwave operation we adiabatically ramp down tweezers to a depth of .
Preserving the coherence of a quantum system as it is scaled up is a known challenge across platforms for quantum computing and simulation kjaergaard_superconducting_2020. This difficulty persists even for neutral atoms, albeit at a lower level, due to residual interactions with a noisy and inhomogeneous electromagnetic environment, particularly with tweezers themselves (due to induced scattering and dephasing). Using far-off-resonant optical tweezers helps to preserve coherence, since the differential light shift of the hyperfine qubit decreases as at constant trap depth, and the scattering rate as , where is the tweezer laser detuning relative to the dominant electronic transition kuhr_analysis_2005, ozeri_hyperfine_2005, ozeri_errors_2007. We verify this by measuring a depolarization time of s (Ext. Data Fig. 7d), and an ensemble dephasing time in the atom array of ms (Fig. 4b), limited by trap depth inhomogeneity. Measured site-by-site, the dephasing time is ms, consistent with being limited by an atomic temperature of during microwave operation kuhr_analysis_2005. In Ext. Data Fig. 8a and in the Methods we present and discuss site-resolved qubit coherence data.
The remaining reversible dephasing can be further mitigated by dynamical decoupling. By applying cycles of XY16 sequences gullion_new_1990, souza_robust_2012 with a period of 12.5 ms between pulses, the measured dephasing time is s, thereby setting a new benchmark for the coherence time of an array of hyperfine qubits in optical tweezers bluvstein_quantum_2022, graham_multi-qubit_2022 (Fig. 4c). In addition, we investigate in Ext. Data Fig. 7g the coherence time at different trap depths, yielding notably s at the full trap depth of mK. Although significantly lower, this result also surpasses previous known results with hyperfine qubits in a tweezer array. In the next section, we leverage the long coherence times demonstrated here to realize high-fidelity qubit transport beugnon_two-dimensional_2007 — the linchpin of zone-based state-of-the-art neutral atom quantum computation bluvstein_quantum_2022, bluvstein_logical_2024, finkelstein_universal_2024 — across record distances.
Finally, we demonstrate global coherent single-qubit operation of our atom array and probe single-qubit gate fidelities through global randomized benchmarking knill_randomized_2008, xia_randomized_2015, nikolov_randomized_2023. To compensate for the inhomogeneous Rabi frequency across the array, which would result in pulse-length errors, we use the SCROFULOUS composite pulse cummins_tackling_2003. We apply gates sampled from the 24 unitaries composing the Clifford group , followed by an inverse operation, and measure the final population in (Fig. 4d). Fitting the decay as the number of gates increases yields an average Clifford gate fidelity , limited by phase noise in our system likely due to magnetic field noise (Methods). This could be readily addressed by upgrading the current sources driving the magnetic field coils, shielding the vacuum cell, or by operating at MHz-scale by driving optical Raman transitions levine_dispersive_2022.
Towards quantum computing with a zone-based architecture — coherent transport
We now focus more specifically on the practical implementation of a quantum computer, as it is a flagship application of our work and because it demands the most sophisticated toolbox of aforementioned use cases. Universal quantum computing requires single-qubit and two-qubit gates which can be implemented either through single-site addressing graham_multi-qubit_2022 or a zone-based architecture bluvstein_quantum_2022, bluvstein_logical_2024.
Zone-based architectures leverage the ability to dynamically move atoms in a coherence-preserving manner beugnon_two-dimensional_2007, bluvstein_quantum_2022, enabling long-range, non-local connectivity, which allows for less stringent quantum error correction bounds than grid-like connectivity bravyi_tradeoffs_2010, presenting a promising pathway to large-scale fault-tolerant quantum computation. A zone-based architecture also provides a path for mid-circuit readout bluvstein_logical_2024.
We depict a possible zone layout in Fig. 5a, which includes a storage zone large enough for more than 6,100 atoms. We do not foresee challenges in creating the zones themselves, e.g., Rydberg-based two-qubit gates should be feasible in a large interaction zone for more than 500 gates in parallel with state-of-the-art fidelities (SI §IV).
However, coherence-preserving transport between storage and adjacent interaction or readout zones might require covering large distances of 500 m. While moving atoms using acousto-optic deflectors (AODs) is now a well-established practice to resort them into a deterministic configuration barredo_atom-by-atom_2016, endres_atom-by-atom_2016 or to transport them coherently bluvstein_quantum_2022, shaw_multi-ensemble_2024, bluvstein_logical_2024, finkelstein_universal_2024, this distance is several times further than previously demonstrated single-atom transport bluvstein_quantum_2022, bluvstein_logical_2024. To address this challenge, we investigate the feasibility of coherent transport at this scale by loading atoms directly into tweezers made by AODs and performing the movement, given that pick-up and drop-off effects are not dependent on the distance of transport. We utilize the most wide-band commercially available AODs at near-infrared wavelengths, which for the optical parameters used here cover up to m along one axis (Methods).
In probing the survival of atoms as they are moved long-distance with differing speeds, we find that the speed of transport is strongly constrained by cylindrical lensing, an effect that occurs as the AOD frequency is rapidly swept dickson_optical_1972, and which is more deleterious as the AOD field of view is scaled up (SI). In Fig. 5b, we present data that illustrate a striking consequence of this effect: with the same trap depth and trajectory, atoms can be transported considerably faster in a diagonal direction, utilizing two crossed AODs, when compared with using one AOD only. Such diagonal moves effectively turn cylindrical lensing into spherical lensing, thereby suppressing the worst impact. In the SI, we describe how diagonal motion can be used to transport atoms between zones without any penalty compared with a usual configuration relying on straight moves. We also discuss other known optical strategies that could completely remove any lensing friedman_acousto-optic_2000.
With diagonal moves, we first demonstrate in Fig. 5c the coherence of atoms being held in static tweezers in comparison with the coherence of atoms being transported by 610 m in 1.6 ms. We suppress any dephasing with one XY4 dynamical decoupling sequence per move, and do not observe loss in coherence between the static and transported case.
However, for any realistic application of coherent transport, atoms will have to be moved multiple times while maintaining coherence. Thus, we characterize the fidelity of the quantum channel defined by the coherent transport, through interleaved randomized benchmarking magesan_efficient_2012 (Fig. 5d). To our knowledge, this the first such rigorous transport channel fidelity characterization for neutral atoms. To maximize the dephasing cancellation, we apply dynamical decoupling in a transformed Clifford frame (Methods).
We perform this benchmarking technique for an average move distance from the storage zone to the interaction zone of m (Fig. 5e,g), and also for a long move distance of m (Fig. 5f). For each set of parameters (Fig. 5e-g), we first characterize the survival probability of an atom in a tweezer at the end of the sequence for different move durations (top panel). For selected durations, we then characterize the actual randomized benchmarking signal, the return probability to the initial quantum state (middle panel).
Notably, the resulting return probability data do not follow an exponential trend, because at large numbers of moves they become dominated by losses from the trap; as such, the fidelity for the transport channel depends on the number of previously executed moves. We thus define an instantaneous fidelity, the fidelity of the transport channel after a certain number of previous moves, shown in the bottom panel (Methods).
The instantaneous fidelity approaches a constant value for small number of moves (), where losses are the sub-dominant error. This regime is most relevant for QEC, since one can in principle swap data qubits and ancilla qubits every few layers of gates chow_circuit-based_2024. We find that the fidelities in this regime approach 99.95% for both transport distances of 270 m (Fig. 5e) and 610 m (Fig. 5g) in ms and in ms respectively using mK deep traps. We note that there is a small discrepancy between the early-time instantaneous fidelities obtained for 270 m and 610 m, and anticipate it could be remedied through more careful trap depth calibration along the trajectory.
For some applications, one might desire to optimize on the speed of movement and use a deeper trap to do so. In Fig. 5g, we show that atoms can be moved by 270 m in 400 s with a trap depth of mK, at the cost of a reduced fidelity of 99.8%. Comparing Fig. 5e and Fig. 5g illustrates an important trade-off pertaining to coherent transport: while atoms can be moved faster by increasing the trap depth , the associated transport fidelity for small number of moves is also reduced. In the limit where noise is entirely induced by tweezer intensity fluctuations, this can be understood by noticing that the dephasing strength scales as when the required duration for long-distance transport merely scales as . We note that experimentally in static traps, we find an even stronger scaling of coherence time than linear in , likely due to other sources of noise (Ext. Data Fig. 7g).
Finally, we envision utilizing multiple pairs of crossed AODs to cover the full extent of the array in Fig. 5a, with the demonstrated long-distance transport allowing overlap between adjacent AOD-pair controlled regions. With the layout presented here, four such regions would be necessary. Alternatively, additional scanning techniques, e.g., fast-scanning mirrors, can be used to position the field of view of single pair of crossed AODs across the full array in an iterative fashion.
As we lay out in detail in the SI, such techniques can be used also for initial rearrangement of atoms into the storage zone. For example, by implementing the recently developed “Tetris algorithm” wang_accelerating_2023 in four quadrants, with conservative estimates for relavant timings based on simulation, data, and previous experiments (SI Table I), we expect that we can sort the array in parallel in 137 ms or sequentially quadrant-by-quadrant in 522 ms. Further, this estimate is based on straight moves along a single AOD axis, which could be sped up by utilizing techniques for mitigating cylindrical lensing (SI §III.B).
Conclusion and outlook
We have shown that while scaling qubit numbers in optical tweezers to more than 6,100, we can simultaneously achieve record-high imaging survival alongside competitive imaging fidelity. We measure record room-temperature vacuum-limited lifetime and coherence time in alkali-metal atom tweezer arrays, with a high global single-qubit gate fidelity that is limited by technical noise. We additionally characterize the fidelity of a long-distance quantum transport channel utilizing randomized benchmarking and providing clear evidence for the potential to use this large scale array with a zone-based architecture for quantum computing (SI).
Our results usher in a new generation of neutral atom quantum processors based on several thousand qubits, as described in the previous section. Additionally, large-scale programmable devices enabling advances in quantum metrology rosenband_exponential_2013, ludlow_optical_2015, norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020, finkelstein_universal_2024 and simulation haferkamp_closing_2020, orourke_entanglement_2023, julia-farre_amorphous_2024 are made accessible through this work. For example, our platform — with the demonstrated qubit numbers — could be used for verifiable quantum advantage with low-depth evolution haferkamp_closing_2020, ringbauer_verifiable_2023. Tweezer clocks could be scaled using near-infrared, high-power tweezers for loading and imaging tao_high-fidelity_2024 before transferring atoms to magic-wavelength traps for clock operation norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020, finkelstein_universal_2024. We also foresee applications in quantum simulation for problems where boundary effects play an important role bernien_probing_2017, browaeys_many-body_2020, ebadi_quantum_2021, scholl_quantum_2021, orourke_entanglement_2023, which can be minimized with the large system sizes demonstrated here.
Finally, our work indicates that further scaling of the optical tweezer array platform to tens of thousands of trapped atoms should be achievable with current technology, while essentially preserving high-fidelity control. In our present apparatus, several factors limit the number of sites. One limitation is the finite number of pixels of each SLM (which reduces the diffraction efficiency as the array size is increased), along with reduced SLM diffraction efficiency at higher incident laser powers. By using available higher-resolution SLMs, and by exploring techniques with higher pixel modulation depth moreno_diffraction_2020, we hope to utilize both power and field of view more efficiently.
Furthermore, we observe worsening optical aberrations at tweezer powers greater than that in the present study due to thermal heating of the objective. This is the main limitation on atom number for the results in this work, even after aberrations were mitigated using the SLM (Methods). This constraint could be circumvented by utilizing an objective with a housing material that retains less heat or with integrated cooling strategies. Such upgrades should allow us to almost double the number of tweezers that we create using two fiber amplifiers. We further anticipate the potential to switch from polarization combination to wavelength-based array combination, opening up further avenues for increasing tweezer number with similar techniques to those utilized in this work (SI). Atom numbers may further be increased in our array with the same number of tweezers by utilizing enhanced loading brown_gray-molasses_2019 or re-loading techniques shaw_dark-state_2023, norcia_iterative_2024_mod. Already in the near term, we expect to increase the number of atomic qubits to over ten thousand with the current system using a subset of these techniques.
Acknowledgements.
We acknowledge insightful discussions with, and feedback from, Adam Shaw, Harry Levine, Zunqi Li, Ran Finkelstein, Pascal Scholl, Joonhee Choi, Dolev Bluvstein, and Soonwon Choi. We acknowledge support from the Gordon and Betty Moore Foundation (Grant GBMF11562), the Weston Havens Foundation, the Institute for Quantum Information and Matter, an NSF Physics Frontiers Center (NSF Grant PHY-2317110), the NSF QLCI program (2016245), the NSF CAREER award (1753386), the Army Research Office MURI program (W911NF2010136), the U.S. Department of Energy (DE-SC0021951), the DARPA ONISQ program (W911NF2010021), and the Air Force Office for Scientific Research Young Investigator Program (FA9550-19-1-0044). Support is also acknowledged from the U.S. Department of Energy, Office of Science, National Quantum Information Science Research Centers, Quantum Systems Accelerator. H.J.M. acknowledges support from the NSF Graduate Research Fellowship Program under Grant No. 2139433. K.H.L. acknowledges support from the AWS-Quantum postdoctoral fellowship and the NUS Development Grant AY2023/2024.Data availability
The data and codes that support the findings of this study are available from the corresponding author upon request.
Author contributions
H.J.M. and M.E. conceived the idea and experiment. H.J.M., G.N., E.B., K.H.L., and X.L. performed numerical simulations and contributed to the experimental setup. H.J.M., G.N., E.B., and K.H.L. performed the experiments. H.J.M., G.N., and E.B. analyzed the data. All authors contributed to the manuscript. M.E. supervised this project.
Competing interests
The authors declare no competing interests.

### v4

@@
 Abstract
-Optical tweezer arrays browaeys_many-body_2020, kaufman_quantum_2021 have had a transformative impact on atomic and molecular physics over the past years, and they now form the backbone for a wide range of leading experiments in quantum computing saffman_quantum_2016, henriet_quantum_2020, bluvstein_quantum_2022, graham_multi-qubit_2022, ma_universal_2022, finkelstein_universal_2024, simulation bernien_probing_2017, browaeys_many-body_2020, ebadi_quantum_2021, scholl_quantum_2021, shaw_benchmarking_2024, and metrology norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020.
-Underlying this development is the simplicity of single-particle control and detection inherent to the technique barredo_atom-by-atom_2016, endres_atom-by-atom_2016, kim_situ_2016, liu_building_2018, anderegg_optical_2019.
-Typical experiments trap tens to hundreds of atomic qubits young_half-minute-scale_2020, scholl_quantum_2021, huft_simple_2022, singh_mid-circuit_2023, tao_high-fidelity_2024, bluvstein_logical_2024, shaw_benchmarking_2024, and very recently systems with around one thousand atoms were realized without defining qubits or demonstrating coherent control pause_supercharged_2024, norcia_iterative_2024_mod, gyger_continuous_2024.
-However, scaling to thousands of atomic qubits with long coherence times and low-loss, high-fidelity imaging is an outstanding challenge and critical for progress in quantum computing, simulation, and metrology, in particular, towards applications with quantum error correction preskill_quantum_2018, xu_constant-overhead_2024.
-Here, we experimentally realize an array of optical tweezers trapping over 6,100 neutral atoms in around 12,000 sites while simultaneously surpassing state-of-the-art performance for several critical metrics that underpin the success of the platform.
-Specifically, while scaling to such a large number of atoms, we also demonstrate a coherence time of 12.6(1) seconds, a record for hyperfine qubits in an optical tweezer array bluvstein_quantum_2022, graham_multi-qubit_2022.
-Further, we show trapping lifetimes close to 23 minutes in a room-temperature apparatus, enabling record-high imaging survival covey_2000-times_2019, blodgett_imaging_parsed_2023 of in combination with an imaging fidelity of over .
-We lay out a detailed, near-term plan to enable zone-based quantum computing bluvstein_quantum_2022, bluvstein_logical_2024 with atoms, and demonstrate a crucial ingredient thereof: coherent moves of up to m with a fidelity of , as characterized through interleaved randomized benchmarking.
-Our results, together with other recent developments evered_high-fidelity_2023, ma_high-fidelity_2023, zhang_scaled_2024, finkelstein_universal_2024, indicate that universal quantum computing bluvstein_quantum_2022, graham_multi-qubit_2022, ma_universal_2022, finkelstein_universal_2024 with ten thousand atomic qubits could be a near-term prospect.
-Furthermore, our work could pave the way for quantum simulation orourke_entanglement_2023, julia-farre_amorphous_2024 and metrology ludlow_optical_2015, rosenband_exponential_2013, haferkamp_closing_2020 experiments with inherent single particle readout and positioning capabilities at a similar scale.
-Optical tweezer arrays browaeys_many-body_2020, kaufman_quantum_2021 have transformed atomic and molecular physics experiments by simplifying detection and enabling control at the individual-particle level barredo_atom-by-atom_2016, endres_atom-by-atom_2016, kim_situ_2016, liu_building_2018, anderegg_optical_2019, resulting in rapid, recent progress in quantum computing saffman_quantum_2016, henriet_quantum_2020, bluvstein_quantum_2022, graham_multi-qubit_2022, ma_universal_2022, finkelstein_universal_2024, quantum simulation bernien_probing_2017, browaeys_many-body_2020, ebadi_quantum_2021, scholl_quantum_2021, shaw_benchmarking_2024, and optical frequency metrology norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020.
-They are part of broader efforts towards single atom control in optical traps, including lattices miroshnychenko_atom-sorting_2006, weitenberg_single-spin_2011 and lattice-tweezer hybrid systems young_half-minute-scale_2020, tao_high-fidelity_2024, gyger_continuous_2024.
-For quantum computers, simulators of spin systems and tweezer clocks, each atom typically encodes a single qubit that is controlled with electromagnetic fields, and ideally features long coherence times to enable these applications with high fidelity.
-Such optically trapped atomic qubits coexist with other platforms that have single-qubit control and readout, including ion traps bruzewicz_trapped-ion_2019 and superconducting qubits kjaergaard_superconducting_2020.
+Optical tweezer arrays browaeys_many-body_2020, kaufman_quantum_2021 have transformed atomic and molecular physics, now forming the backbone for a range of leading experiments in quantum computing saffman_quantum_2016, henriet_quantum_2020, bluvstein_quantum_2022, graham_multi-qubit_2022, ma_universal_2022, finkelstein_universal_2024, simulation bernien_probing_2017, browaeys_many-body_2020, ebadi_quantum_2021, scholl_quantum_2021, shaw_benchmarking_2024, and metrology norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020.
+Typical experiments trap tens to hundreds of atomic qubits young_half-minute-scale_2020, scholl_quantum_2021, huft_simple_2022, singh_mid-circuit_2023, tao_high-fidelity_2024, bluvstein_logical_2024, shaw_benchmarking_2024, and recently systems with around one thousand atoms were realized without defining qubits or demonstrating coherent control pause_supercharged_2024, norcia_iterative_2024_mod, gyger_continuous_2024.
+However, scaling to thousands of atomic qubits with long coherence times, low-loss, and high-fidelity imaging is an outstanding challenge and critical for progress in quantum science, particularly towards quantum error correction preskill_quantum_2018, xu_constant-overhead_2024.
+Here, we experimentally realize an array of optical tweezers trapping over 6,100 neutral atoms in around 12,000 sites, simultaneously surpassing state-of-the-art performance for several metrics that underpin the success of the platform.
+Specifically, while scaling to such a large number of atoms, we demonstrate a coherence time of 12.6(1) seconds, a record for hyperfine qubits in an optical tweezer array bluvstein_quantum_2022, graham_multi-qubit_2022.
+We show room-temperature trapping lifetimes of 23 minutes, enabling record-high imaging survival covey_2000-times_2019, blodgett_imaging_parsed_2023 of with an imaging fidelity of over .
+We present a plan for zone-based quantum computing bluvstein_quantum_2022, bluvstein_logical_2024 and demonstrate necessary coherence-preserving qubit transport and pick-up/drop-off operations on large spatial scales, characterized through interleaved randomized benchmarking.
+Our results, along with recent developments evered_high-fidelity_2023, ma_high-fidelity_2023, zhang_scaled_2024, finkelstein_universal_2024, indicate that universal quantum computing bluvstein_quantum_2022, graham_multi-qubit_2022, ma_universal_2022, finkelstein_universal_2024 and quantum error correction preskill_quantum_2018, xu_constant-overhead_2024 with thousands to tens of thousands of physical qubits could be a near-term prospect.
+Optical tweezer arrays browaeys_many-body_2020, kaufman_quantum_2021 have transformed atomic and molecular physics experiments by simplifying detection and enabling individual-particle control barredo_atom-by-atom_2016, endres_atom-by-atom_2016, kim_situ_2016, liu_building_2018, anderegg_optical_2019, resulting in rapid, recent progress in quantum computing saffman_quantum_2016, henriet_quantum_2020, bluvstein_quantum_2022, graham_multi-qubit_2022, ma_universal_2022, finkelstein_universal_2024, simulation bernien_probing_2017, browaeys_many-body_2020, ebadi_quantum_2021, scholl_quantum_2021, shaw_benchmarking_2024, and metrology norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020.
+In this context, each atom typically encodes a single qubit that is controlled with electromagnetic fields, and ideally features long coherence times to enable these applications with high fidelity.
+Such optically trapped atomic qubits, including lattices miroshnychenko_atom-sorting_2006, weitenberg_single-spin_2011 and lattice-tweezer hybrid systems young_half-minute-scale_2020, tao_high-fidelity_2024, gyger_continuous_2024, coexist with other platforms that have single-qubit control and readout, including ion traps bruzewicz_trapped-ion_2019 and superconducting qubits kjaergaard_superconducting_2020.
 There are important incentives to scale up such fully programmable qubit platforms.
-For example, optical clocks gain in stability with increasing atom number ludlow_optical_2015, rosenband_exponential_2013, and proposals have been made for quantum simulation experiments that benefit from several thousand qubits to explore emergent collective behavior orourke_entanglement_2023, julia-farre_amorphous_2024 or to demonstrate verifiable quantum advantage haferkamp_closing_2020, ringbauer_verifiable_2023.
-In particular, quantum error correction (QEC) requires large system sizes, while maintaining high fidelities and long coherence times, in order to scale up the number of logical qubits preskill_quantum_2018, google_quantum_ai_suppressing_2023.
-Even the most resource-efficient QEC protocols require at least several thousand physical qubits with an error-per-gate level of to encode 100 error-corrected logical qubits bravyi_high-threshold_2024, xu_constant-overhead_2024.
-However, published results from universal quantum computing architectures based on ion traps and superconducting qubits currently leverage only tens to hundreds of qubits moses_race_2023, kim_evidence_2023, and with current technologies for fabrication and control, generally suffer from deleterious effects as they are scaled kjaergaard_superconducting_2020, bruzewicz_trapped-ion_2019.
-Neutral atoms in optical tweezer arrays are a promising platform towards rapid scalability in the near term thanks to a programmable architecture that is readily adaptable to larger system sizes.
-In particular, universal quantum computing has recently been realized in such systems bluvstein_quantum_2022, graham_multi-qubit_2022, ma_universal_2022, finkelstein_universal_2024, based on demonstrations of individual qubit addressing huie_repetitive_2023, lis_midcircuit_2023, norcia_midcircuit_2023, shaw_multi-ensemble_2024, high-fidelity entangling gates evered_high-fidelity_2023, finkelstein_universal_2024, and coherence-preserving dynamical reconfigurability beugnon_two-dimensional_2007, bluvstein_quantum_2022, finkelstein_universal_2024, alongside ancilla-based mid-circuit measurement singh_mid-circuit_2023, bluvstein_logical_2024, finkelstein_universal_2024 and erasure error detection scholl_erasure_2023, ma_high-fidelity_2023.
-Leveraging the versatility of the optical tweezer architecture, several applications have been shown, including executing quantum phase estimation graham_multi-qubit_2022, realizing the toric code structure bluvstein_quantum_2022, building an optical clock with ancilla-based readout finkelstein_universal_2024, and implementing logical qubit operations bluvstein_logical_2024.
-In terms of current system sizes, tens to hundreds of qubits are controlled by typical tweezer array experiments young_half-minute-scale_2020, scholl_quantum_2021, huft_simple_2022, singh_mid-circuit_2023, tao_high-fidelity_2024, bluvstein_logical_2024, shaw_benchmarking_2024.
-Very recently, tweezer systems with about a thousand atoms have been realized in a discontiguous array based on interleaved microlens elements pause_supercharged_2024, and via repeated reloading from a reservoir norcia_iterative_2024_mod, gyger_continuous_2024, following earlier iterative reloading schemes singh_dual-element_2022, shaw_dark-state_2023; none of these experiments, however, report control of qubits or measurement of coherence times.
-Here, we demonstrate a tweezer array with 11,998 sites, that traps over 6,100 atomic qubits, simultaneously surpassing state-of-the-art values for metrics underpinning the usefulness of the platform, including hyperfine qubit coherence time, trapping lifetime in a room-temperature apparatus, coherent transport distance and fidelity, as well as combined imaging fidelity and survival (Fig.
+Optical atomic clocks gain stability with increasing atom number ludlow_optical_2015, rosenband_exponential_2013, while quantum simulation experiments benefit from thousands of qubits to explore emergent collective phenomena orourke_entanglement_2023, julia-farre_amorphous_2024 or demonstrate verifiable quantum advantage haferkamp_closing_2020, ringbauer_verifiable_2025.
+Most critically, quantum error correction (QEC) demands both large system sizes and exceptional fidelities: even the most resource-efficient protocols require several thousand physical qubits operating with error rates below to encode 100 logical qubits bravyi_high-threshold_2024, xu_constant-overhead_2024.
+This represents a fundamental scalability challenge that has limited the practical impact of quantum technologies.
+Current universal quantum computing architectures, such as those based on neutral atoms bluvstein_quantum_2022, graham_multi-qubit_2022, ma_universal_2022, finkelstein_universal_2024, ions moses_race_2023, and superconducting qubits kim_evidence_2023, typically operate with tens to hundreds of qubits.
+While most platforms suffer from increasingly deleterious effects as system size grows kjaergaard_superconducting_2020, bruzewicz_trapped-ion_2019, neutral atoms in optical tweezer arrays offer a promising solution for rapid scalability in the near term thanks to a programmable architecture adaptable to larger system sizes.
+Universal quantum computing capabilities with neutral atoms have recently been realized in optical tweezer array systems, based on demonstrations of individual qubit addressing huie_repetitive_2023, lis_midcircuit_2023, norcia_midcircuit_2023, shaw_multi-ensemble_2024, high-fidelity entangling gates evered_high-fidelity_2023, finkelstein_universal_2024, coherence-preserving dynamical reconfigurability beugnon_two-dimensional_2007, bluvstein_quantum_2022, finkelstein_universal_2024, and ancilla-based mid-circuit measurement singh_mid-circuit_2023, bluvstein_logical_2024, finkelstein_universal_2024.
+Very recently, tweezer systems with about a thousand atoms have been realized in a discontiguous array based on interleaved microlens elements pause_supercharged_2024, and via repeated reloading from a reservoir norcia_iterative_2024_mod, gyger_continuous_2024; none of these experiments, however, report control of qubits, measurement of coherence times, or coherence-preserving transport.
+Here, we demonstrate a tweezer array with 11,998 sites, that traps over 6,100 atomic qubits, simultaneously matching or surpassing state-of-the-art values for metrics underpinning the usefulness of the platform, including hyperfine qubit coherence time, trapping lifetime in a room-temperature apparatus, coherent transport distance and fidelity, trap transfer fidelity, as well as combined imaging fidelity and survival (Fig.
 1).
-Our results have implications for aforementioned applications in quantum science, in particular, concerning large-scale quantum computing and error correction, as discussed in more detail below.
+Our results have implications for the aforementioned applications in quantum science, in particular, concerning large-scale quantum computing and error correction, as discussed in more detail below.
 Summary of approach and results
-Our approach leverages high-power trapping of single atoms at far-off-resonant wavelengths in a specially designed, room-temperature vacuum chamber (Methods, Ext.
+Our approach leverages high-power trapping of single cesium-133 atoms at far-off-resonant wavelengths in a specially designed, room-temperature vacuum chamber (Methods, Ext.
 Data Fig.
@@
 1e).
-Specifically, we demonstrate imaging of single cesium-133 atoms with a survival probability of , in combination with an imaging fidelity of , surpassing the state-of-the-art achieved in much smaller arrayscovey_2000-times_2019.
-This, alongside a 22.9(1) minute vacuum-limited lifetime in our room-temperature apparatus schymik_single_2021 — significantly longer than typical state-of-the-art vacuum lifetimes for tweezer arrays in room-temperature apparatuses— provides realistic timescales for array operations in large scale arrays with minimal loss, e.g., for atomic rearrangement barredo_atom-by-atom_2016, endres_atom-by-atom_2016, kim_situ_2016.
+Specifically, we demonstrate single-atom imaging with a survival probability of and a fidelity of , surpassing the state-of-the-art achieved in much smaller arrayscovey_2000-times_2019.
+This, alongside a 22.9(1)-minute vacuum-limited lifetime in our room-temperature apparatus schymik_single_2021—significantly longer than typical state-of-the-art vacuum lifetimes for tweezer arrays in room-temperature apparatuses—provides realistic timescales for array operations in large scale arrays with minimal loss; e.g., for atomic rearrangement barredo_atom-by-atom_2016, endres_atom-by-atom_2016, kim_situ_2016.
 Importantly, we further demonstrate a coherence time of 12.6(1) s, a record for a hyperfine qubit tweezer array, surpassing previous values by almost an order of magnitude bluvstein_quantum_2022, graham_multi-qubit_2022 and approaching results for a single hyperfine qubit in a customized blue-detuned trap tian_extending_2024, alkali atoms in an optical lattice wu_sterngerlach_2019, and nuclear qubits in a tweezer array barnes_assembly_2022.
-We also show a single-qubit gate fidelity of measured with global randomized benchmarking, limited by technical noise.
-We observe high uniformity across the sites in the array for loading probability (Ext.
-Data Fig.
-2c), trap depth (Ext.
-Data Fig.
-2d), imaging survival (Ext.
-Data Fig.
-5a), scattering rate (Ext.
-Data Fig.
-4d), and qubit frequency (Ext.
-Data Fig.
-8a).
-Finally, we lay out a realistic plan to implement zone-based universal quantum computation bluvstein_quantum_2022, bluvstein_logical_2024 with 6,000 atoms in the near term.
-We summarize in this work the key elements of our proposal, elucidated further in the Supplementary Information (SI).
-We notably demonstrate coherent atomic transport on a distance of 610 m with a fidelity of 99.95%, a crucial ingredient for scaling this architecture.
+We also show a single-qubit gate fidelity of measured with global randomized benchmarking.
+Finally, we demonstrate coherent atomic transport across 610 m with a fidelity of 99.95% and coherent transfer between static and dynamic traps with a fidelity of 99.81(3)%.
+These together form crucial ingredients for scaling atomic quantum processors in a zone-based architecture, with a detailed plan laid out further in the Supplementary Information (SI).
 Our results indicate that quantum computing with 6,000 atomic qubits is a near-term prospect, providing a path towards QEC with hundreds of logical qubits xu_constant-overhead_2024.
 Large-scale optical tweezer generation
-We scale the optical tweezer array platform to 11,998 tweezers, while prioritizing achieving long qubit coherence times and high fidelities (Fig.
-1e).
-As such, we trap using near-infrared wavelengths, far-detuned from dominant electric-dipole transitions to minimize hyperfine qubit decoherence from photon scattering and dephasing processes ozeri_hyperfine_2005, ozeri_errors_2007.
-Cesium atoms possess the highest polarizability among the stable alkali-metal atoms at near-infrared wavelengths where commercial fiber amplifiers routinely provide continuous-wave laser powers that exceed 100 W.
-Thus, a large number of traps can be generated with sufficient depth.
+To scale the optical tweezer array platform, while extending hyperfine coherence times, we generate traps using near-infrared wavelengths, far-detuned from dominant electric-dipole transitions, thus minimizing photon scattering and dephasing processes ozeri_hyperfine_2005, ozeri_errors_2007.
+Cesium atoms possess the highest polarizability among the stable alkali-metal atoms at near-infrared wavelengths where commercial fiber amplifiers provide continuous-wave laser powers that exceed 100 W.
+Thus, a large number of traps can be created with sufficient depth.
 A representative single shot image of the array is shown in Fig.
@@
 1b.
-The atoms are spaced by and held in traps at 1055 nm and 1061 nm, generated using spatial light modulators (SLMs), whose hologram phases are optimized with a weighted Gerchberg-Saxton algorithm kim_gerchberg-saxton_2019, kim_large-scale_2019 to uniformize the tweezer trap depth across the array (Methods).
-The tweezer light is combined with polarization and focused through a high numerical aperture objective with an unusually large field of view of 1.5 mm diameter that provides a large area for atom trapping and manipulation in ultra-high vacuum (Fig.
+The atoms are spaced by and held in traps at 1055 nm and 1061 nm, generated using spatial light modulators (SLMs), whose hologram phases are optimized with a weighted Gerchberg-Saxton algorithm kim_gerchberg-saxton_2019, kim_large-scale_2019 to uniformize the tweezer trap depth (Methods).
+The tweezer light is combined with polarization and focused through a high numerical aperture objective with a large field of view of 1.5 mm in diameter, usable for atom trapping and manipulation (Fig.
 1c).
+The tweezers are created with 130 W of optical power generated from fiber amplifiers.
+After transmission through the optical path, around 35-40 W reaches the objective, and from trap parameter measurements (see “Tweezer generation” Methods section), we estimate mW to be used per tweezer at the atom plane.
 We measure an average trap depth of mK, with a standard deviation of across all sites (Ext.
@@
 2d), enabling consistent loading probability per site.
-Using Raman sideband spectroscopy with a 9.2 GHz amplitude-modulation technique levine_dispersive_2022, we measure radial and axial trapping frequencies of 29.3 kHz and 5.6 kHz, respectively (for the traps generated with the 1061 nm laser only; see Ext.
-Data Fig.
-9).
-The setup used to address the atomic motion with a Raman configuration is detailed in the Methods.
-From this measurement we infer a tweezer waist .
 Loading and imaging single atoms
@@
 Before imaging the atoms, we use a multi-pronged approach to filter out atoms in spurious off-plane traps, residual from the SLM tweezer creation (Methods).
-We then zero the magnetic field and apply two-dimensional polarization gradient cooling (2D PGC) beams that are approximately parallel to the atom array plane (- plane in Fig.
-1c) for fluorescence imaging of single atoms, which also simultaneously cool the atoms.
-The photons scattered from each atom are collected on a quantitative CMOS (qCMOS) camera for 80 ms, using the same objective with which the traps are created.
-We find that each site is loaded with a single atom on average 51.2% of the time with a relative standard deviation of 3.4% across the sites, demonstrating uniform filling of single atoms in a large tweezer array (Ext.
+We then zero the magnetic field and apply two-dimensional polarization gradient cooling (2D PGC) in the atom array plane (- plane in Fig.
+1c) for fluorescence imaging of single atoms, which simultaneously cools the atoms.
+Imaging light is applied for 80 ms, and photons are imaged on a quantitative CMOS camera (qCMOS).
+We find that each site has a loading probability of 51.2% with a relative standard deviation of 3.4% across the sites, demonstrating uniform filling of single atoms (Ext.
 Data Fig.
@@
 1d).
-We then detect and distinguish atomic presence in the array with high fidelity.
-Each image undergoes a binarization procedure (detailed in Methods) whereby each site is attributed a value of 0 (no atom detected) or 1 (one atom detected).
-Note that we detect no more than one atom schlosser_sub-poissonian_2001 in each tweezer.
+We distinguish atomic presence in the array with high fidelity.
+Each image undergoes a binarization procedure (Methods) whereby each site is attributed a value of 0 (no atom detected) or 1 (one atom detected).
 We weight the collected photons in a -pixel box centered around each site cooper_alkaline-earth_2018, so as to add more weight to pixels close to the center of each site’s point-spread function (Ext.
@@
 From this we find an imaging fidelity of 99.99374(8)% (note that we excise the first image, which we find has slightly lower fidelity and survival probability; Methods).
-Crucial to this result are the homogeneous photon scattering rate across the array (Ext.
+Crucial to this result is the homogeneous photon scattering rate across the array (Ext.
 Data Fig.
-4d) and the consistency of the point-spread function even for sites at the edge of the array (waist radius of 1.7 pixels with a standard deviation of 0.2 pixels).
-Furthermore, we find that treating each site separately with an individual threshold only marginally improves the imaging fidelity to 99.9939(1)%, indicating that the imaging parameters are sufficiently consistent across the atom array.
+4d) and the consistency of the point-spread function across the array (waist radius of 1.7 pixels with a standard deviation of 0.2 pixels).
+Consistent imaging parameters across the array further are evidenced in that we find that treating each site with an individual threshold only marginally improves the imaging fidelity to 99.9939(1)%.
 Finally, we estimate that the imaging fidelity in the absence of atomic loss would be closer to 99.999% (Methods).
 Imaging survival and vacuum-limited lifetime
-The probability of not losing a single atom in a tweezer array during imaging and due to finite vacuum lifetime both decrease exponentially in the number of atoms in an array, making these crucial metrics to optimize for large-scale array operation.
-The vacuum-limited lifetime, in particular, sets the ultimate upper bound on the duration during which operations can be executed without loss of an atom in a given experimental setup.
-This can, for example, be applied to give an upper limit on the fidelity with which one can achieve a defect-free array via atom rearrangement barredo_atom-by-atom_2016, endres_atom-by-atom_2016.
-We probe the vacuum-limited lifetime by using an empirically optimized cooling sequence consisting of a 10 ms 2D PGC cooling block every 2 s.
+The probability of losing no atom in a tweezer array during imaging and due to finite vacuum lifetime both decrease exponentially in the number of atoms in an array, making these crucial metrics to optimize for large-scale array operation.
+The vacuum-limited lifetime, in particular, sets an upper bound on the duration during which operations can be executed without loss of an atom in a given experimental run.
+This can, for example, be applied as an upper limit on the fidelity with which one can achieve a defect-free array via atom rearrangement barredo_atom-by-atom_2016, endres_atom-by-atom_2016.
+We probe the vacuum-limited lifetime using an empirically optimized cooling sequence consisting of a 10-ms 2D PGC cooling block every 2 s.
 By fitting the exponential decay of the atom survival, we find a lifetime of 22.9(1) min (Fig.
 3a).
-This is a significantly long timescale compared with state-of-the-art room-temperature atomic experiments, and only within a factor of five of the longest reported lifetime in a cryogenic apparatus schymik_single_2021.
+This is a significantly long timescale compared with state-of-the-art room-temperature atomic experiments, and within a factor of five of the longest reported lifetime in a cryogenic apparatus schymik_single_2021.
 The result indicates that the probability of losing a single atom across the entire array remains under 50% during 100 ms, a relevant timescale for dynamical array reconfiguration and quantum processor operation.
-Moreover, we accurately characterize the imaging survival probability, without assuming any parameters, by performing 80 ms repeated imaging up to 1,000 times, after which % of initially loaded atoms still survive (Fig.
+Moreover, we accurately characterize the imaging survival probability, without assuming any parameters, by performing 80-ms repeated imaging up to 1,000 times, after which % of initially loaded atoms still survive (Fig.
 3b).
@@
 In cesium atoms, the hyperfine ground states ( and ) provide such a subspace for storing quantum information (see Methods for state preparation and readout procedures).
-Furthermore, entanglement via Rydberg interactions can be readily transferred to this qubit to realize high-fidelity two-qubit gates evered_high-fidelity_2023, finkelstein_universal_2024.
-We now demonstrate the storage and manipulation of quantum information in a large-scale atom array by measuring the coherence time and global single-qubit gate fidelity.
-To this end we use a microwave horn to drive the hyperfine transition (Fig.
+Furthermore, entanglement via Rydberg interactions can be readily transferred to this qubit to realize high-fidelity two-qubit gates evered_high-fidelity_2023.
+We demonstrate the storage and manipulation of quantum information in a large-scale atom array by measuring the coherence time and global single-qubit gate fidelity using a microwave horn to drive the hyperfine transition (Fig.
 4a).
@@
 Preserving the coherence of a quantum system as it is scaled up is a known challenge across platforms for quantum computing and simulation kjaergaard_superconducting_2020.
-This difficulty persists even for neutral atoms, albeit at a lower level, due to residual interactions with a noisy and inhomogeneous electromagnetic environment, particularly with tweezers themselves (due to induced scattering and dephasing).
-Using far-off-resonant optical tweezers helps to preserve coherence, since the differential light shift of the hyperfine qubit decreases as at constant trap depth, and the scattering rate as , where is the tweezer laser detuning relative to the dominant electronic transition kuhr_analysis_2005, ozeri_hyperfine_2005, ozeri_errors_2007.
-We verify this by measuring a depolarization time of s (Ext.
+This difficulty persists even for neutral atoms, albeit at a lower level, due to residual interactions with a noisy and inhomogeneous electromagnetic environment, particularly with the tweezers themselves.
+Thus, we choose to trap in far-off-resonant optical tweezers to preserve coherence, since at constant trap depth the differential light shift of the hyperfine qubit decreases as and the scattering rate as , where is the tweezer laser detuning relative to the dominant electronic transition kuhr_analysis_2005, ozeri_hyperfine_2005, ozeri_errors_2007.
+We indeed observe long coherence times, measuring a depolarization time of s (Ext.
 Data Fig.
-7d), and an ensemble dephasing time in the atom array of ms (Fig.
+7d), and an array-averaged ensemble dephasing time of ms (Fig.
 4b), limited by trap depth inhomogeneity.
@@
 Data Fig.
-8a and in the Methods we present and discuss site-resolved qubit coherence data.
-The remaining reversible dephasing can be further mitigated by dynamical decoupling.
-By applying cycles of XY16 sequences gullion_new_1990, souza_robust_2012 with a period of 12.5 ms between pulses, the measured dephasing time is s, thereby setting a new benchmark for the coherence time of an array of hyperfine qubits in optical tweezers bluvstein_quantum_2022, graham_multi-qubit_2022 (Fig.
+8 and in the Methods we present and discuss site-resolved qubit coherence data.
+The dephasing can be further mitigated by dynamical decoupling.
+By applying cycles of XY16 sequences gullion_new_1990 with a period of 12.5 ms between pulses, the measured dephasing time is s, a new benchmark for the coherence time of an array of hyperfine qubits in optical tweezers bluvstein_quantum_2022, graham_multi-qubit_2022 (Fig.
 4c).
@@
 7g the coherence time at different trap depths, yielding notably s at the full trap depth of mK.
-Although significantly lower, this result also surpasses previous known results with hyperfine qubits in a tweezer array.
-In the next section, we leverage the long coherence times demonstrated here to realize high-fidelity qubit transport beugnon_two-dimensional_2007 — the linchpin of zone-based state-of-the-art neutral atom quantum computation bluvstein_quantum_2022, bluvstein_logical_2024, finkelstein_universal_2024 — across record distances.
-Finally, we demonstrate global coherent single-qubit operation of our atom array and probe single-qubit gate fidelities through global randomized benchmarking knill_randomized_2008, xia_randomized_2015, nikolov_randomized_2023.
-To compensate for the inhomogeneous Rabi frequency across the array, which would result in pulse-length errors, we use the SCROFULOUS composite pulse cummins_tackling_2003.
-We apply gates sampled from the 24 unitaries composing the Clifford group , followed by an inverse operation, and measure the final population in (Fig.
+Although lower, this result also surpasses previous known results with hyperfine qubits in a tweezer array.
+Finally, we determine single-qubit gate fidelities through global randomized benchmarking knill_randomized_2008, xia_randomized_2015, nikolov_randomized_2023.
+To compensate for the inhomogeneous Rabi frequency across the array, we use the SCROFULOUS composite pulse cummins_tackling_2003.
+We apply gates sampled from the Clifford group , followed by an inverse operation, and measure the final population in (Fig.
 4d).
 Fitting the decay as the number of gates increases yields an average Clifford gate fidelity , limited by phase noise in our system likely due to magnetic field noise (Methods).
-This could be readily addressed by upgrading the current sources driving the magnetic field coils, shielding the vacuum cell, or by operating at MHz-scale by driving optical Raman transitions levine_dispersive_2022.
-Towards quantum computing with a zone-based architecture — coherent transport
+This could be readily addressed by upgrading the current sources driving the magnetic field coils or by operating at MHz-scale Rabi rates with optical Raman transitions levine_dispersive_2022.
+Towards zone-based quantum computing — coherent long-distance transport and atom transfer
 We now focus more specifically on the practical implementation of a quantum computer, as it is a flagship application of our work and because it demands the most sophisticated toolbox of aforementioned use cases.
-Universal quantum computing requires single-qubit and two-qubit gates which can be implemented either through single-site addressing graham_multi-qubit_2022 or a zone-based architecture bluvstein_quantum_2022, bluvstein_logical_2024.
-Zone-based architectures leverage the ability to dynamically move atoms in a coherence-preserving manner beugnon_two-dimensional_2007, bluvstein_quantum_2022, enabling long-range, non-local connectivity, which allows for less stringent quantum error correction bounds than grid-like connectivity bravyi_tradeoffs_2010, presenting a promising pathway to large-scale fault-tolerant quantum computation.
-A zone-based architecture also provides a path for mid-circuit readout bluvstein_logical_2024.
-We depict a possible zone layout in Fig.
-5a, which includes a storage zone large enough for more than 6,100 atoms.
+Universal quantum computing requires single-qubit and two-qubit gates which have been implemented either through single-site addressing graham_multi-qubit_2022 or a zone-based architecture bluvstein_quantum_2022, bluvstein_logical_2024.
+Zone-based architectures leverage the ability to dynamically move atoms in a coherence-preserving manner beugnon_two-dimensional_2007, bluvstein_quantum_2022, enabling long-range, non-local connectivity, which allows for less stringent quantum error correction bounds bravyi_tradeoffs_2010.
+This architecture also provides a path for mid-circuit readout bluvstein_logical_2024.
+We depict a possible zone layout in Ext.
+Data Fig.
+10a and the Supplementary Information (SI), which includes a storage zone large enough for more than 6,100 atoms.
 We do not foresee challenges in creating the zones themselves, e.g., Rydberg-based two-qubit gates should be feasible in a large interaction zone for more than 500 gates in parallel with state-of-the-art fidelities (SI §IV).
 However, coherence-preserving transport between storage and adjacent interaction or readout zones might require covering large distances of 500 m.
-While moving atoms using acousto-optic deflectors (AODs) is now a well-established practice to resort them into a deterministic configuration barredo_atom-by-atom_2016, endres_atom-by-atom_2016 or to transport them coherently bluvstein_quantum_2022, shaw_multi-ensemble_2024, bluvstein_logical_2024, finkelstein_universal_2024, this distance is several times further than previously demonstrated single-atom transport bluvstein_quantum_2022, bluvstein_logical_2024.
-To address this challenge, we investigate the feasibility of coherent transport at this scale by loading atoms directly into tweezers made by AODs and performing the movement, given that pick-up and drop-off effects are not dependent on the distance of transport.
-We utilize the most wide-band commercially available AODs at near-infrared wavelengths, which for the optical parameters used here cover up to m along one axis (Methods).
-In probing the survival of atoms as they are moved long-distance with differing speeds, we find that the speed of transport is strongly constrained by cylindrical lensing, an effect that occurs as the AOD frequency is rapidly swept dickson_optical_1972, and which is more deleterious as the AOD field of view is scaled up (SI).
-In Fig.
-5b, we present data that illustrate a striking consequence of this effect: with the same trap depth and trajectory, atoms can be transported considerably faster in a diagonal direction, utilizing two crossed AODs, when compared with using one AOD only.
-Such diagonal moves effectively turn cylindrical lensing into spherical lensing, thereby suppressing the worst impact.
-In the SI, we describe how diagonal motion can be used to transport atoms between zones without any penalty compared with a usual configuration relying on straight moves.
-We also discuss other known optical strategies that could completely remove any lensing friedman_acousto-optic_2000.
+While moving atoms using acousto-optic deflectors (AODs) is now a well-established practice to resort them into a deterministic configuration barredo_atom-by-atom_2016, endres_atom-by-atom_2016 or to transport them coherently bluvstein_quantum_2022, shaw_multi-ensemble_2024, bluvstein_logical_2024, finkelstein_universal_2024, this distance is significantly further than previously demonstrated distances for single-atom transport with tweezers bluvstein_quantum_2022, bluvstein_logical_2024.
+Furthermore, transferring atoms between dynamic (AOD-generated) and static (SLM-generated) tweezers requires precise relative alignment, conceivably challenging in our system due to the high laser power or potential for worsening aberrations over the large FOV.
+Thus, we investigate the feasibility of coherence-preserving transport and SLM-AOD trap transfer over larger length scales.
+First, isolating challenges with the coherent transport operation, we load atoms directly into ten AOD-generated tweezers and characterize coherent moves up to 610 m (Fig.
+5, top section of Ext.
+Data Fig.
+10).
+Second, we assess the viability of large-scale parallel AOD-SLM trap transfers with 195 AOD-generated tweezers that span a square of 504 m 468 m (Fig.
+6).
+As an outlook, we demonstrate a proof-of-principle combination of these techniques in a large-scale static array (although in a different trap layout) moving a 2D array of 47 atoms over 375 m, a distance comparable to predicted zone spacings in our system (Ext.
+Data Fig.
+10e, f).
+For all operations, we utilize the most wide-band commercially available AODs at near-infrared wavelengths, which cover up to m along one axis for the optical parameters used here (Methods).
+In probing atom survival as a function of long-distance movement speeds, we find that the speed of transport is strongly constrained by cylindrical lensing—an effect that occurs when the AOD frequency is rapidly swept dickson_optical_1972—which becomes increasingly deleterious as the AOD field of view is increased (SI III.A).
+Notably, using a pair of crossed AODs for diagonal transport converts cylindrical lensing into spherical lensing, enabling significantly faster movement (Fig.
+5a).
 With diagonal moves, we first demonstrate in Fig.
-5c the coherence of atoms being held in static tweezers in comparison with the coherence of atoms being transported by 610 m in 1.6 ms.
-We suppress any dephasing with one XY4 dynamical decoupling sequence per move, and do not observe loss in coherence between the static and transported case.
-However, for any realistic application of coherent transport, atoms will have to be moved multiple times while maintaining coherence.
-Thus, we characterize the fidelity of the quantum channel defined by the coherent transport, through interleaved randomized benchmarking magesan_efficient_2012 (Fig.
-5d).
-To our knowledge, this the first such rigorous transport channel fidelity characterization for neutral atoms.
+5b negligible loss of coherence for atoms transported by 610 m in 1.6 ms.
+We suppress dephasing with one XY4 dynamical decoupling sequence per move.
+Realistic applications of coherent transport involve multiple consecutive moves.
+Therefore, we characterize the fidelity of the quantum channel defined by coherent transport through interleaved randomized benchmarking magesan_efficient_2012 (IRB, Fig.
+5c, Methods).
+To the best of our knowledge, such a quantitative characterization of transport fidelity in neutral atoms has not been previously demonstrated.
 To maximize the dephasing cancellation, we apply dynamical decoupling in a transformed Clifford frame (Methods).
-We perform this benchmarking technique for an average move distance from the storage zone to the interaction zone of m (Fig.
-5e,g), and also for a long move distance of m (Fig.
-5f).
-For each set of parameters (Fig.
-5e-g), we first characterize the survival probability of an atom in a tweezer at the end of the sequence for different move durations (top panel).
-For selected durations, we then characterize the actual randomized benchmarking signal, the return probability to the initial quantum state (middle panel).
-Notably, the resulting return probability data do not follow an exponential trend, because at large numbers of moves they become dominated by losses from the trap; as such, the fidelity for the transport channel depends on the number of previously executed moves.
-We thus define an instantaneous fidelity, the fidelity of the transport channel after a certain number of previous moves, shown in the bottom panel (Methods).
-The instantaneous fidelity approaches a constant value for small number of moves (), where losses are the sub-dominant error.
-This regime is most relevant for QEC, since one can in principle swap data qubits and ancilla qubits every few layers of gates chow_circuit-based_2024.
-We find that the fidelities in this regime approach 99.95% for both transport distances of 270 m (Fig.
-5e) and 610 m (Fig.
-5g) in ms and in ms respectively using mK deep traps.
-We note that there is a small discrepancy between the early-time instantaneous fidelities obtained for 270 m and 610 m, and anticipate it could be remedied through more careful trap depth calibration along the trajectory.
-For some applications, one might desire to optimize on the speed of movement and use a deeper trap to do so.
-In Fig.
-5g, we show that atoms can be moved by 270 m in 400 s with a trap depth of mK, at the cost of a reduced fidelity of 99.8%.
-Comparing Fig.
-5e and Fig.
-5g illustrates an important trade-off pertaining to coherent transport: while atoms can be moved faster by increasing the trap depth , the associated transport fidelity for small number of moves is also reduced.
-In the limit where noise is entirely induced by tweezer intensity fluctuations, this can be understood by noticing that the dephasing strength scales as when the required duration for long-distance transport merely scales as .
-We note that experimentally in static traps, we find an even stronger scaling of coherence time than linear in , likely due to other sources of noise (Ext.
+We perform this benchmarking technique for a distance of m (Fig.
+5d), with diagonal moves.
+We first measure the survival probability of an atom in a tweezer at the end of the sequence for different move durations (top panel).
+For a 1.6-ms move using mK-deep traps, we then characterize the return probability to the initial quantum state after IRB as a function of the number of moves (middle panel).
+Other distances, trap depths and move times are shown in Ext.
 Data Fig.
-7g).
-Finally, we envision utilizing multiple pairs of crossed AODs to cover the full extent of the array in Fig.
-5a, with the demonstrated long-distance transport allowing overlap between adjacent AOD-pair controlled regions.
-With the layout presented here, four such regions would be necessary.
-Alternatively, additional scanning techniques, e.g., fast-scanning mirrors, can be used to position the field of view of single pair of crossed AODs across the full array in an iterative fashion.
-As we lay out in detail in the SI, such techniques can be used also for initial rearrangement of atoms into the storage zone.
-For example, by implementing the recently developed “Tetris algorithm” wang_accelerating_2023 in four quadrants, with conservative estimates for relavant timings based on simulation, data, and previous experiments (SI Table I), we expect that we can sort the array in parallel in 137 ms or sequentially quadrant-by-quadrant in 522 ms.
-Further, this estimate is based on straight moves along a single AOD axis, which could be sped up by utilizing techniques for mitigating cylindrical lensing (SI §III.B).
+10.
+The resulting IRB return probability data is non-exponential in the number of moves, because at large numbers of moves, trap losses become dominant and the fidelity for the transport channel depends on the number of previously executed moves.
+This motivates defining an instantaneous fidelity; i.e., the fidelity of the transport channel after a certain number of previous moves (Methods), shown in the bottom panel of Fig.
+5d.
+The instantaneous fidelity approaches a constant value of for small numbers of one-way moves (), where losses are the sub-dominant error.
+This regime is most relevant for QEC, since data qubits and ancilla qubits can, in principle, be swapped every few layers of gates chow_circuit-based_2024_new.
+We then move on to characterizing the atom transfer between static and dynamic tweezers.
+We demonstrate that these operations proceed without the emergence of unexpected technical challenges by performing high-fidelity parallel AOD-SLM transfer across the full field of view of the AOD (Fig.
+6).
+We use 195 AOD tweezers spread across 504 m 468 m (Fig.
+6a) to perform and characterize the repeated transfer procedure, post-selected on initially filled SLM sites.
+As with coherent transport benchmarking, we evaluate the transfer fidelity as a function of the number of one-way transfers through IRB (Fig.
+6d).
+To execute faster (or higher-fidelity transfers at a given duration), we propose and implement a trajectory where AOD ramp-and-move operations are simultaneously optimized with machine learning techniques to maximize survival (Fig.
+6c, Methods).
+Compared with our manually optimized trajectory (Fig.
+6b), this technique yields significantly higher atomic survival, and enables a one-way transfer fidelity of 99.81(3)% for transfers (compare blue and gray data in Fig.
+6d, top panel).
+In the future, such machine learning techniques could also be used to optimize combined pick-up and transport, where we find a fidelity of for the first operations at the chosen timescales with manually optimized methods (Methods, Ext.
+Data Fig.
+10f).
+Finally, to cover the full extent of the array, we envision utilizing multiple pairs of crossed AODs, with the demonstrated long-distance transport allowing overlap between adjacent AOD-pair controlled regions (SI Fig.
+2).
+With the layout presented in Ext.
+Data Fig.
+10a and the SI, four such regions would be necessary.
+Alternatively, additional scanning techniques (e.g., fast-scanning mirrors) can be used to position the field of view of a single pair of crossed AODs across the full array iteratively.
+Such techniques are also applicable to initial rearrangement of atoms in the storage zone.
+For example, by implementing a parallel assembly algorithmtian_parallel_2023, wang_accelerating_2023 in four quadrants (SI §II), with estimates for relevant timings based on simulation, data, and previous experiments (SI Table I), we expect that we can sort the array in parallel in 137 ms or sequentially quadrant-by-quadrant in 522 ms.
 Conclusion and outlook
-We have shown that while scaling qubit numbers in optical tweezers to more than 6,100, we can simultaneously achieve record-high imaging survival alongside competitive imaging fidelity.
-We measure record room-temperature vacuum-limited lifetime and coherence time in alkali-metal atom tweezer arrays, with a high global single-qubit gate fidelity that is limited by technical noise.
-We additionally characterize the fidelity of a long-distance quantum transport channel utilizing randomized benchmarking and providing clear evidence for the potential to use this large scale array with a zone-based architecture for quantum computing (SI).
-Our results usher in a new generation of neutral atom quantum processors based on several thousand qubits, as described in the previous section.
+We have shown scaling of neutral atom qubit numbers in optical tweezers to more than 6,100.
+We simultaneously achieve high imaging survival and fidelity as well as a long room-temperature vacuum-limited lifetime.
+We find record coherence times in alkali-metal atom tweezer arrays and a high global single-qubit gate fidelity, limited by technical noise.
+Further, we additionally characterize the fidelity of quantum transport channels for moves and trap transfer at relevant length scales, utilizing randomized benchmarking.
+Our results usher in a new generation of neutral atom quantum processors based on several thousand qubits, particularly relevant for QEC bravyi_high-threshold_2024, xu_constant-overhead_2024.
 Additionally, large-scale programmable devices enabling advances in quantum metrology rosenband_exponential_2013, ludlow_optical_2015, norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020, finkelstein_universal_2024 and simulation haferkamp_closing_2020, orourke_entanglement_2023, julia-farre_amorphous_2024 are made accessible through this work.
-For example, our platform — with the demonstrated qubit numbers — could be used for verifiable quantum advantage with low-depth evolution haferkamp_closing_2020, ringbauer_verifiable_2023.
+For example, our platform—with the demonstrated qubit numbers—could be used for verifiable quantum advantage with low-depth evolution haferkamp_closing_2020, ringbauer_verifiable_2025.
 Tweezer clocks could be scaled using near-infrared, high-power tweezers for loading and imaging tao_high-fidelity_2024 before transferring atoms to magic-wavelength traps for clock operation norcia_seconds-scale_2019, madjarov_atomic-array_2019, young_half-minute-scale_2020, finkelstein_universal_2024.
@@
 In our present apparatus, several factors limit the number of sites.
-One limitation is the finite number of pixels of each SLM (which reduces the diffraction efficiency as the array size is increased), along with reduced SLM diffraction efficiency at higher incident laser powers.
-By using available higher-resolution SLMs, and by exploring techniques with higher pixel modulation depth moreno_diffraction_2020, we hope to utilize both power and field of view more efficiently.
+One limitation is the finite number of pixels of each SLM (reducing the diffraction efficiency as the array size is increased), along with reduced SLM diffraction efficiency at higher incident laser powers.
+By using available higher resolution SLMs, and by exploring techniques with higher pixel modulation depth moreno_diffraction_2020, we hope to utilize both power and field of view more efficiently.
 Furthermore, we observe worsening optical aberrations at tweezer powers greater than that in the present study due to thermal heating of the objective.
@@
 Such upgrades should allow us to almost double the number of tweezers that we create using two fiber amplifiers.
-We further anticipate the potential to switch from polarization combination to wavelength-based array combination, opening up further avenues for increasing tweezer number with similar techniques to those utilized in this work (SI).
+We further anticipate the potential to switch from polarization combination to wavelength-based array combination, opening up further avenues for increasing tweezer number with similar techniques to those utilized in this work.
 Atom numbers may further be increased in our array with the same number of tweezers by utilizing enhanced loading brown_gray-molasses_2019 or re-loading techniques shaw_dark-state_2023, norcia_iterative_2024_mod.
@@
 We acknowledge support from the Gordon and Betty Moore Foundation (Grant GBMF11562), the Weston Havens Foundation, the Institute for Quantum Information and Matter, an NSF Physics Frontiers Center (NSF Grant PHY-2317110), the NSF QLCI program (2016245), the NSF CAREER award (1753386), the Army Research Office MURI program (W911NF2010136), the U.S.
-Department of Energy (DE-SC0021951), the DARPA ONISQ program (W911NF2010021), and the Air Force Office for Scientific Research Young Investigator Program (FA9550-19-1-0044).
+Department of Energy (DE-SC0021951), the DARPA ONISQ program (W911NF2010021), the Air Force Office for Scientific Research Young Investigator Program (FA9550-19-1-0044), and the Heising-Simons Foundation (2024-4852).
 Support is also acknowledged from the U.S.